OpenStudy (anonymous):

The average value of F(x)= -1/(x^2) on (1/2,1) is ?.

OpenStudy (bahrom7893):

Integral from 1/2 to 1 of f(x)dx time 1/(b-a)

OpenStudy (bahrom7893):

so it will be: 1 / (1-1/2) Integral from 1/2 to 1, etc..

OpenStudy (anonymous):

so do i take the derevative or the anti ?

OpenStudy (bahrom7893):

2 Integral from 1/2 to 1 of -1/(x^2) dx

OpenStudy (anonymous):

oh ok so how i thought when i i need to go to an inegral i take the derevative

OpenStudy (bahrom7893):

no u just integrate..

OpenStudy (anonymous):

ok ty again

OpenStudy (anonymous):

the answer im getting is -6 which is not a choice

OpenStudy (bahrom7893):

-1 * 2 = -2

OpenStudy (bahrom7893):

check the integral

OpenStudy (anonymous):

sorry idk what im doing wrong...2(-1/(1/4))-2(-1/1) so -8+2=-6 but that is clearly not right

OpenStudy (bahrom7893):

Integral of -1/x^2 = Int [-1x^(-2) ] = -1 * x^(-2+1)/(-2+1) = x^(-1) from 1/2 to 1

OpenStudy (bahrom7893):

1^(-1) - (1/2)^(-1) = 1 - 2 = -1

OpenStudy (bahrom7893):

then 2*-1 = -2

OpenStudy (anonymous):

haha tyvm it just clicked lol